OpenStudy (anonymous):

how to graph 3x-6y=12?

OpenStudy (anonymous):

Just rearrange the equation. 3x-6y=12 3x-12=6y 3/6x-12/6=y 1/2x-2=y Go to geogebra and enter the rearranged equation on the input line. http://www.geogebra.org/cms/